OpenStudy (anonymous):

How do you solve this integral? Integral from 0 to 2 |x^2-1| dx

OpenStudy (anonymous):

Can you draw the graph of it? it usually helps

OpenStudy (anonymous):

int(x-a)= int(x)-int(a) feel free to integrate them separately, int from 0 to 2 of int(x^2)-int(1).

OpenStudy (anonymous):

a parabola down 1 unit and the negative part flipped?

OpenStudy (anonymous):

exactly @jojo921 Now, about the negative part, it runs from x=0 to x=1, so you can compute that integral normally and just take the absolute value of it

OpenStudy (anonymous):

Then you maybe want to set up an additional integral that runs from x=1 to x=2

OpenStudy (anonymous):

evaluate((x^3/3)-x) from 0 to 2

OpenStudy (anonymous):

I've tried it that way.. but still received the wrong answer.

OpenStudy (anonymous):

@Lost_in_Translation it's the absolute value, not the regular x^2-1 function

OpenStudy (anonymous):

what do you get @jojo921 ?

OpenStudy (anonymous):

and maybe post the correct answer right away too if you have it please.

OpenStudy (anonymous):

I've got 4/3 but the correct answer is 2

OpenStudy (anonymous):

your absolutely right i didn't catch that, sorry jojo ignore my previous answer.

OpenStudy (anonymous):

Well your integration should become like \[\int\limits_0^2 \left| x^2-1 \right|dx= \left| \int\limits_0^1(x^2-1)dx \right|+ \int\limits_1^2(x^2-1)dx\]

OpenStudy (anonymous):

and yes, that turns out to be 2.

OpenStudy (anonymous):

maybe i did my calculations wrong. let me try it again and thank you!

OpenStudy (anonymous):

you're welcome @jojo921

OpenStudy (anonymous):

yup i got 2

OpenStudy (anonymous):

way to go (-;